# MOSAIC: Masked Objective with Selective Adaptation for In-domain Contrastive Learning

### v1

@@ Abstract
We introduce MOSAIC (Masked Objective with Selective Adaptation for In-domain Contrastive learning), a multi-stage framework for domain adaptation of sentence embedding models that incorporates joint domain-specific masked supervision. Our approach addresses the challenges of adapting large-scale general-domain sentence embedding models to specialized domains. By jointly optimizing masked language modeling (MLM) and contrastive objectives within a unified training pipeline, our method enables effective learning of domain-relevant representations while preserving the robust semantic discrimination properties of the original model. We empirically validate our approach on both high-resource and low-resource domains, achieving improvements up to 13.4% in NDCG@10 (Normalized Discounted Cumulative Gain) over strong general-domain baselines. Comprehensive ablation studies further demonstrate the effectiveness of each component, highlighting the importance of balanced joint supervision and staged adaptation.
@@
Vera Pavlova1,2, Mohammed Makhlouf2 1burevestnik.ai  2rttl.ai 1v@burevestnik.ai,2v@rttl.ai
@@ 1 Introduction
Self-supervised learning has enabled significant progress in natural language processing, with methods like MLM (Devlin et al., 2019; Liu et al., 2020; Conneau et al., 2020; Sanh et al., 2019) and contrastive training (Reimers and Gurevych, 2019; Wu et al., 2020; Liu et al., 2021) driving recent developments. However, these methods are typically explored separately, as effectively combining MLM and contrastive learning remains a significant challenge, since their joint optimization often results in conflicting training signals and suboptimal performance (Gao et al., 2021). Nevertheless, unifying these objectives presents an opportunity to leverage the complementary strengths of token-level (MLM) and sentence-level (contrastive) supervision, while also improving the quality of learned representations by mitigating the anisotropy problem (a phenomenon that confines embeddings to a narrow cone-like region in the vector space, thereby limiting their expressiveness) (Ethayarajh, 2019; Li et al., 2020; Gao et al., 2021). While there have been successful attempts to combine MLM and contrastive objectives for training language models (Meng et al., 2021; Chi et al., 2021) and sentence embeddings (Gao et al., 2021; Wu et al., 2022; Giorgi et al., 2021), the majority of the prior work has focused on general-domain data.
General-domain sentence embedding models are now widely available, many trained on vast general-domain corpora using a two-stage approach: an initial pre-training phase on massive unlabeled data, followed by supervised fine-tuning (Wang et al., 2022b; Li et al., 2023; Nussbaum et al., 2024; Merrick et al., 2024). The data used for pre-training can exceed half a billion sentence pairs (hundreds of gigabytes of text), resources that are rarely available in specific domains. Although these general-domain models can perform competitively in specialized areas, their lack of domain-specific knowledge often limits performance. To address this gap, we propose domain adaptation of pre-trained embedding models that leverage their ability to distinguish between similar and dissimilar pairs and transfer it to a domain-specific embedding model.
Previous research on language model adaptation highlights the importance of domain-specific vocabulary for improving results on downstream tasks (Beltagy et al., 2019; Gu et al., 2020a). However, simply adding domain-specific vocabulary and continuing MLM training degrades the contrastive properties of the learned representations, since the encoder loses its desirable characteristics under the token prediction objective (Wu et al., 2022). On the other hand, adding new tokens and continuing only with the contrastive objective provides insufficient training signals to update new domain tokens, as the embedding matrix receives diluted signals due to the pooling functions applied to generate sentence embeddings.
@@
Figure 1: The MOSAIC pipeline: a multi-stage framework for domain adaptation of sentence embedding models.
To thoroughly evaluate our method, we apply it to both high-resource and low-resource domains. Most domain-adaptation research focuses on high-resource and medium-resource domains, which is valuable for benchmarking, comparison with strong baselines, and conducting ablation studies. Yet this focus restricts the generalizability of adaptation methods to truly low-resource domains, which are common in real-world applications. Such domains often face acute data scarcity, making robust adaptation methods essential for ensuring equitable access to state-of-the-art language technologies and maximizing the real-world impact of embedding models. To demonstrate the robustness and practical value of our approach, we validate it in two domains: the Biomedical domain, which is characterized by high-resource scientific texts, and the Islamic domain, which represents low-resource but culturally significant content. This allows us to test the robustness of our method even when there is very limited in-domain data.
Our main contributions are as follows: (1) We propose MOSAIC, a novel domain adaptation approach for pretrained sentence embedding models that jointly optimizes MLM and contrastive objectives within a mutual information maximization framework. (2) We empirically validate our method on both high-resource (biomedical) and low-resource (Islamic) domains, demonstrating substantial gains over strong general-domain baselines. (3) We conduct comprehensive ablation studies to analyze the contribution of each component and the dynamics of joint objective training. (4) We release our code and pretrained models to support reproducibility and facilitate future research.
@@ 2.1 Contrastive and MLM Objectives
Prior work has explored joint contrastive and masked objectives in two distinct directions: large-scale language model pretraining and general-purpose sentence embedding models. COCO-LM integrates contrastive learning into a pretraining pipeline for transformer language models (Vaswani et al., 2017; Meng et al., 2021), replacing BERT’s Next Sentence Prediction (NSP) objective with a more effective contrastive signal by pairing corrupted and truncated versions of a sentence. The model jointly learns to align these pairs via contrastive loss and to correct the corruption through token-level denoising. COCO-LM demonstrated consistent gains on GLUE tasks (Wang et al., 2018), showing that contrastive objectives outperform NSP for general-purpose pretraining. InfoXLM reframes masked language modeling as a contrastive prediction task, formulating it with the InfoNCE loss (Chi et al., 2021) and applying it to multilingual language model pretraining. Combined with a sentence-level cross-lingual contrastive objective, this joint training enables InfoXLM to achieve state-of-the-art results on cross-lingual understanding and retrieval benchmarks.
For training sentence embeddings, DeCLUTR explicitly combines MLM with contrastive training for sentence embeddings (Giorgi et al., 2021). They construct positive pairs from contiguous spans of the same document, apply BERT-style masking to the anchor span, and train jointly through a single encoder. DeCLUTR was evaluated using SentEval, which largely consists of classification tasks along with some STS and NLI tasks. The authors of InfoCSE introduced a sophisticated framework (Wu et al., 2022) and, rather than combining MLM and contrastive loss on the same encoder output, employed an auxiliary lightweight encoder. While InfoCSE demonstrated gains on STS benchmarks, its architectural separation constrains the transfer of token-level supervision into sentence-level embeddings, limiting the potential benefits of joint training.
Distinct from these methods, our work introduces a joint MLM-contrastive objective explicitly designed for domain adaptation of existing strong sentence embedding models, which, to the best of our knowledge, represents a novel and previously unexplored direction.
@@ 2.2 Domain Adaptation
Domain adaptation is most commonly performed at the language modeling stage, where general-purpose models undergo continued pre-training on in-domain corpora (Lee et al., 2019; Alsentzer et al., 2019). Such approaches typically suffer from the absence of domain-specific vocabulary, which often necessitates training from scratch (Beltagy et al., 2019; Gu et al., 2020a). To avoid these GPU-heavy methods, recent work has explored lightweight domain adaptation by introducing new domain vocabulary to already well-trained models, thereby expediting the pre-training process (Poerner et al., 2020; Sachidananda et al., 2021; Pavlova and Makhlouf, 2023).
For sentence embedding models that have already undergone contrastive training, prior efforts have either relied on small adapters (Huang et al., 2023; Schopf et al., 2023) which often suffer from limited expressiveness, or on data-driven methods such as augmentation, denoising objectives, or generative pseudo-labeling (Thakur et al., 2021; Wang et al., 2021, 2022a). In contrast, we introduce a full model-driven approach to domain adaptation that operates after language model pretraining. By modifying the training objective itself, our method provides a more expressive and principled mechanism for adapting sentence embeddings.
@@ 3.1 Augmenting Contrastive Models with Domain-Specific Vocabulary (MOSAIC-Stage1)
To leverage the robust encoder learned during contrastive pretraining, we reuse both the encoder and the original embedding matrix. However, to accommodate a word distribution shift from a general domain vocabulary to a new domain vocabulary, we augment the model with new domain-specific tokens (see Figure  1):
@@ 3.3 Contrastive-Only Training (MOSAIC-Stage3)
For the third stage, we continue training our model using only the contrastive objective, after the new domain tokens have been introduced and learned. This stage serves as a corrective step, allowing the encoder to recover and reinforce sentence-level discrimination, which may be diluted during joint MLM+contrastive training. By focusing solely on contrastive learning, the model re-aligns its representations to produce robust sentence embeddings.
@@
nomic-embed-text-v1 86.471 41.48 66.648 37.0082 34.3009 35.028 70.500 79.923 56.670
@@ 4 Experiments on a High-Resource Domain
Fine-tuning data. We evaluate our models in a zero-shot setting. To prevent benchmark data leakage (a common concern for sentence embedding models), we fine-tune on BioASQ Task 9a (Tsatsaronis et al., 2015), which does not overlap with any of the MTEB evaluation datasets. The dataset consists of approximately 16 million PubMed title–abstract pairs annotated with MeSH (Medical Subject Headings); Importantly, this fine-tuning step is not part of the proposed MOSAIC training pipeline. Instead, it corresponds to preparing the final supervised model (MOSAIC-Biosuper) used for evaluation in supervise setting.
@@
Baselines. For unsupervised baselines, we use nomic-embed-text-v1unsup [^1] https://huggingface.co/nomic-ai/nomic-embed-text-v1-unsupervised as our primary baseline representing an unsupervised contrastive embedding model pretrained on general-domain data. We also train this model on the unsupervised training data described above and include the nomic-embed-bio model in the comparison. To analyze the impact of each stage on domain adaptation, we evaluate three variants of our MOSAIC pipeline, all initialized from the nomic-embed-text-v1unsup checkpoint. MOSAIC-Stage1 augments the model with new domain-specific vocabulary without further pretraining. MOSAIC-Stage2 applies joint training with masked language modeling and contrastive objectives on domain data. MOSAIC-Stage3 continues training with the contrastive objective alone, yielding the final unsupervised MOSAIC model. For supervised baselines, we select a diverse set of well-established embedding models that report MTEB scores on biomedical datasets, as listed on the official MTEB leaderboard [^2] https://huggingface.co/spaces/mteb/leaderboard. As noted in Section 2, most prior efforts on sentence embedding adaptation have focused on data-driven techniques and do not produce a general-purpose model checkpoint suitable for evaluation or reuse. Moreover, the biomedical embedding landscape remains sparse: to the best of our knowledge, there is no research-based, dedicated biomedical sentence embedding model available on MTEB. To address this gap, we prepare a biomedical sentence embedding model derived from the state-of-the-art PubMedBERT, which we further train on the domain-specific data described above (PubMedBERT-embed).
@@
Implementation Details. We implement the joint MLM and contrastive training on top of the Nomic repository [^3] https://github.com/nomic-ai/contrastors/tree/main. For the purely contrastive stage, we reuse the original implementation from the repository. The model architecture is based on BERT (Devlin et al., 2019) with several modifications introduced by the Nomic repo. In MOSAIC pipeline, we first expand the tokenizer by adding approximately 9k new biomedical tokens, and then implement a joint training objective that combines masked language modeling with contrastive learning. We set the masking rate to 0.15, the MLM loss weighting hyperparameter $\alpha=0.3$ throughout the joint training phase. Details of ablation on $\alpha$ and masking rate can be found in Section 4.2, and other hyperparameter settings are provided in Appendix E. We train stages two and three of the proposed pipeline using only in-batch negatives, and additionally include hard-mined negatives during fine-tuning.
4.1 Results
@@
Our results demonstrate several important trends regarding domain adaptation for sentence embeddings (see Table 1). First, we observe that simply continuing pretraining a general domain embedding model (nomic-embed-text-v1unsup (Nussbaum et al., 2024)) on in-domain data can lead to reduced performance compared to the original general-domain baseline across most benchmarks (as in the nomic-embed-bio model), suggesting that naive in-domain adaptation may distort learned representations. This issue becomes even more pronounced when augmenting the vocabulary with domain-specific tokens without any retraining (MOSAIC-Bio-Stage1), resulting in a substantial performance drop across all datasets, likely due to embedding mismatch. In contrast, applying a joint MLM+contrastive objective (MOSAIC-Bio-Stage2) recovers and further improves results, while a final contrastive-only training stage (MOSAIC-Bio-Stage3) achieves the highest scores on four benchmarks (BiorxivClusteringP2P, MedicalQARetrieval, MedrxivClusteringP2P, and SciFact), resulting in the best average performance overall and with a 2.8% increase over the general-domain baseline. These results highlight a clear progression across adaptation stages, where naive vocabulary expansion leads to degradation, targeted joint supervision restores model quality, and a final contrastive stage enables robust domain adaptation. In the supervised setting, our MOSAIC-Biosuper achieves the highest average score and the best scores on BIOSSES, MedrxivClusteringP2P, and MedicalQARetrieval, outperforming other strong supervised baselines on these key biomedical tasks (additional analysis on results on TRECCOVID can be found in Appendix C).

### v2

@@ Abstract
We introduce MOSAIC (Masked Objective with Selective Adaptation for In-domain Contrastive learning), a multi-stage framework for domain adaptation of text embedding models that incorporates joint domain-specific masked supervision. Our approach addresses the challenges of adapting large-scale general-domain text embedding models to specialized domains. By jointly optimizing masked language modeling (MLM) and contrastive objectives within a unified training pipeline, our method enables effective learning of domain-relevant representations while preserving the robust semantic discrimination properties of the original model. We empirically validate our approach on both high-resource and low-resource domains, achieving improvements up to 13.4% in NDCG@10 (Normalized Discounted Cumulative Gain) over strong general-domain baselines. Comprehensive ablation studies further demonstrate the effectiveness of each component, highlighting the importance of balanced joint supervision and staged adaptation.
@@
Vera Pavlova1,2, Mohammed Makhlouf2 1burevestnik.ai  2rttl.ai 1v@burevestnik.ai,2v@rttl.ai, mm@rttl.ai
@@ 1 Introduction
Self-supervised learning has enabled significant progress in natural language processing, with methods like MLM (Devlin et al., 2019; Liu et al., 2020; Conneau et al., 2020; Sanh et al., 2019) and contrastive training (Reimers and Gurevych, 2019; Wu et al., 2020; Liu et al., 2021) driving recent developments. However, these methods are typically explored separately, as effectively combining MLM and contrastive learning remains a significant challenge, since their joint optimization often results in conflicting training signals and suboptimal performance (Gao et al., 2021). Nevertheless, unifying these objectives presents an opportunity to leverage the complementary strengths of token-level (MLM) and sentence-level (contrastive) supervision, while also improving the quality of learned representations by mitigating the anisotropy problem (a phenomenon that confines embeddings to a narrow cone-like region in the vector space, thereby limiting their expressiveness) (Ethayarajh, 2019; Li et al., 2020; Gao et al., 2021). While there have been successful attempts to combine MLM and contrastive objectives for training language models (Meng et al., 2021; Chi et al., 2021) and text embeddings (Gao et al., 2021; Wu et al., 2022; Giorgi et al., 2021), the majority of the prior work has focused on general-domain data.
General-domain text embedding models are now widely available, many trained on vast general-domain corpora using a two-stage approach: an initial pre-training phase on massive unlabeled data, followed by supervised fine-tuning (Wang et al., 2022b; Li et al., 2023; Nussbaum et al., 2024; Merrick et al., 2024). The data used for pre-training can exceed half a billion sentence pairs (hundreds of gigabytes of text), resources that are rarely available in specific domains. Although these general-domain models can perform competitively in specialized areas, their lack of domain-specific knowledge often limits performance. To address this gap, we propose domain adaptation of pre-trained embedding models that leverage their ability to distinguish between similar and dissimilar pairs and transfer it to a domain-specific embedding model.
Previous research on language model adaptation highlights the importance of domain-specific vocabulary for improving results on downstream tasks (Beltagy et al., 2019; Gu et al., 2020a). However, simply adding domain-specific vocabulary and continuing MLM training degrades the contrastive properties of the learned representations, since the encoder loses its desirable characteristics under the token prediction objective (Wu et al., 2022). On the other hand, adding new tokens and continuing only with the contrastive objective provides insufficient training signals to update new domain tokens, as the embedding matrix receives diluted signals due to the pooling functions applied to generate text embeddings.
@@
Figure 1: The MOSAIC pipeline: a multi-stage framework for domain adaptation of text embedding models.
To thoroughly evaluate our method, we apply it to both high-resource and low-resource domains. Most domain-adaptation research focuses on high-resource and medium-resource domains, which is valuable for benchmarking, comparison with strong baselines, and conducting ablation studies. Yet this focus restricts the generalizability of adaptation methods to truly low-resource domains, which are common in real-world applications. Such domains often face acute data scarcity, making robust adaptation methods essential for ensuring equitable access to state-of-the-art language technologies and maximizing the real-world impact of embedding models (Pavlova and Makhlouf, 2025). To demonstrate the robustness and practical value of our approach, we validate it in two domains: the Biomedical domain, which is characterized by high-resource scientific texts, and the Islamic domain, which represents low-resource but culturally significant content. This allows us to test the robustness of our method even when there is very limited in-domain data.
Our main contributions are as follows: (1) We propose MOSAIC, a novel domain adaptation approach for pretrained text embedding models that jointly optimizes MLM and contrastive objectives within a mutual information maximization framework. (2) We empirically validate our method on both high-resource (biomedical) and low-resource (Islamic) domains, demonstrating substantial gains over strong general-domain baselines. (3) We conduct comprehensive ablation studies to analyze the contribution of each component and the dynamics of joint objective training. (4) We release our code and pretrained models to support reproducibility and facilitate future research  [^1] https://github.com/rttl-ai/mosaic.
@@ 2.1 Contrastive and MLM Objectives
Prior work has explored joint contrastive and masked objectives in two distinct directions: large-scale language model pretraining and general-purpose text embedding models. COCO-LM integrates contrastive learning into a pretraining pipeline for transformer language models (Vaswani et al., 2017; Meng et al., 2021), replacing BERT’s Next Sentence Prediction (NSP) objective with a more effective contrastive signal by pairing corrupted and truncated versions of a sentence. The model jointly learns to align these pairs via contrastive loss and to correct the corruption through token-level denoising. COCO-LM demonstrated consistent gains on GLUE tasks (Wang et al., 2018), showing that contrastive objectives outperform NSP for general-purpose pretraining. InfoXLM reframes masked language modeling as a contrastive prediction task, formulating it with the InfoNCE loss (Chi et al., 2021) and applying it to multilingual language model pretraining. Combined with a sentence-level cross-lingual contrastive objective, this joint training enables InfoXLM to achieve state-of-the-art results on cross-lingual understanding and retrieval benchmarks.
For training text embeddings, DeCLUTR explicitly combines MLM with contrastive training for text embeddings (Giorgi et al., 2021). They construct positive pairs from contiguous spans of the same document, apply BERT-style masking to the anchor span, and train jointly through a single encoder. DeCLUTR was evaluated using SentEval, which largely consists of classification tasks along with some STS and NLI tasks. The authors of InfoCSE introduced a sophisticated framework (Wu et al., 2022) and, rather than combining MLM and contrastive loss on the same encoder output, employed an auxiliary lightweight encoder. While InfoCSE demonstrated gains on STS benchmarks, its architectural separation constrains the transfer of token-level supervision into sentence-level embeddings, limiting the potential benefits of joint training.
Distinct from these methods, our work introduces a joint MLM-contrastive objective explicitly designed for domain adaptation of existing strong text embedding models, which, to the best of our knowledge, represents a novel and previously unexplored direction.
@@ 2.2 Domain Adaptation
Domain adaptation is most commonly performed at the language modeling stage, where general-purpose models undergo continued pre-training on in-domain corpora (Lee et al., 2019; Alsentzer et al., 2019). Such approaches typically suffer from the absence of domain-specific vocabulary, which often necessitates training from scratch (Beltagy et al., 2019; Gu et al., 2020a). To avoid these GPU-heavy methods, recent work has explored lightweight domain adaptation by introducing new domain vocabulary to already well-trained models, thereby expediting the pre-training process (Poerner et al., 2020; Sachidananda et al., 2021; Pavlova and Makhlouf, 2023; Pavlova, 2023).
For text embedding models that have already undergone contrastive training, prior efforts have either relied on small adapters (Huang et al., 2023; Schopf et al., 2023) which often suffer from limited expressiveness, or on data-driven methods such as augmentation, denoising objectives, or generative pseudo-labeling (Thakur et al., 2021; Wang et al., 2021, 2022a). In contrast, we introduce a full model-driven approach to domain adaptation that operates after language model pretraining. By modifying the training objective itself, our method provides a more expressive and principled mechanism for adapting text embeddings.
@@ 3.1 Augmenting Contrastive Models with Domain-Specific Vocabulary (MOSAIC-Stage1)
To leverage the robust encoder learned during contrastive pretraining, we reuse both the encoder and the original embedding matrix. However, to accommodate a word distribution shift from a general domain vocabulary to a new domain vocabulary, we augment the model with new domain-specific tokens (see Figure  1). The size of the augmented vocabulary in MOSAIC is neither manually selected nor treated as a hyperparameter. Instead, it is a direct and deterministic function of the underlying domain corpus. Given a fixed base tokenizer and a fixed domain corpus, we follow a two-step procedure:
@@ 3.3 Contrastive-Only Training (MOSAIC-Stage3)
For the third stage, we continue training our model using only the contrastive objective, after the new domain tokens have been introduced and learned. This stage serves as a corrective step, allowing the encoder to recover and reinforce sentence-level discrimination, which may be diluted during joint MLM+contrastive training. By focusing solely on contrastive learning, the model re-aligns its representations to produce robust text embeddings.
@@
nomic-embed-text-v1super 86.471 41.48 66.648 37.0082 34.3009 35.028 70.500 79.923 56.670
nomic-embed-biosuper 70.55 35.35 52.19 31.19 25.20 31.37 47.45 56.48 43.72
@@ 4 Experiments on a High-Resource Domain
Fine-tuning data. We evaluate our models in a zero-shot setting. To prevent benchmark data leakage (a common concern for text embedding models), we fine-tune on BioASQ Task 9a (Tsatsaronis et al., 2015), which does not overlap with any of the MTEB evaluation datasets. The dataset consists of approximately 16 million PubMed title–abstract pairs annotated with MeSH (Medical Subject Headings); Importantly, this fine-tuning step is not part of the proposed MOSAIC training pipeline. Instead, it corresponds to preparing the final supervised model (MOSAIC-Biosuper) used for evaluation in supervise setting.
@@
Baselines. For unsupervised baselines, we use nomic-embed-text-v1unsup [^2] https://huggingface.co/nomic-ai/nomic-embed-text-v1-unsupervised as our primary baseline representing an unsupervised contrastive embedding model pretrained on general-domain data. We also train this model on the unsupervised training data described above and include the nomic-embed-bio model in the comparison. To analyze the impact of each stage on domain adaptation, we evaluate three variants of our MOSAIC pipeline, all initialized from the nomic-embed-text-v1unsup checkpoint. MOSAIC-Stage1 augments the model with new domain-specific vocabulary without further pretraining. MOSAIC-Stage2 applies joint training with masked language modeling and contrastive objectives on domain data. MOSAIC-Stage3 continues training with the contrastive objective alone, yielding the final unsupervised MOSAIC model. For supervised baselines, we select a diverse set of well-established embedding models that report MTEB scores on biomedical datasets, as listed on the official MTEB leaderboard [^3] https://huggingface.co/spaces/mteb/leaderboard. As noted in Section 2, most prior efforts on text embedding adaptation have focused on data-driven techniques and do not produce a general-purpose model checkpoint suitable for evaluation or reuse. Moreover, the biomedical embedding landscape remains sparse: to the best of our knowledge, there is no research-based, dedicated biomedical sentence embedding model available on MTEB. To address this gap, we prepare a biomedical text embedding model derived from the state-of-the-art PubMedBERT, which we further train on the domain-specific data described above (PubMedBERT-embed). We additionally include an adapter-based baseline (E5-base-peft-biomed), where LoRA adapters are applied to E5base on the same domain corpus
@@
Implementation Details. We implement the joint MLM and contrastive training on top of the Nomic repository [^4] https://github.com/nomic-ai/contrastors/tree/main. For the purely contrastive stage, we reuse the original implementation from the repository. The model architecture is based on BERT (Devlin et al., 2019) with several modifications introduced by the Nomic repo. In MOSAIC pipeline, we first expand the tokenizer by adding approximately 9k new biomedical tokens, and then implement a joint training objective that combines masked language modeling with contrastive learning. We set the masking rate to 0.15, the MLM loss weighting hyperparameter $\alpha=0.3$ throughout the joint training phase. Details of ablation on $\alpha$ and masking rate can be found in Section 4.2, and other hyperparameter settings are provided in Appendix C. We train stages two and three of the proposed pipeline using only in-batch negatives, and additionally include hard-mined negatives during fine-tuning.
4.1 Results and Analysis
@@
Our results demonstrate several important trends regarding domain adaptation for text embeddings (see Table 1). First, among unsupervised baselines, we observe that simply continuing pretraining a general-domain embedding model (nomic-embed-text-v1unsup) on in-domain data can lead to reduced performance compared to the original general-domain baseline across most benchmarks (as in the nomic-embed-bio model), suggesting that naive in-domain adaptation may distort learned representations. This issue becomes even more pronounced when augmenting the vocabulary with domain-specific tokens without any retraining (MOSAIC-Bio-Stage1), resulting in a substantial performance drop across all datasets, likely due to embedding mismatch. In contrast, applying a joint MLM+contrastive objective (MOSAIC-Bio-Stage2) recovers and further improves results, while a final contrastive-only training stage (MOSAIC-Bio-Stage3) achieves the highest scores on four benchmarks (BiorxivClusteringP2P, MedicalQARetrieval, MedrxivClusteringP2P, and SciFact), resulting in the best average performance overall with a 2.8% increase over the general-domain baseline. These results highlight a clear progression across adaptation stages, where naive vocabulary expansion leads to degradation, targeted joint supervision restores model quality, and a final contrastive stage enables robust domain adaptation.
In the analysis of supervised baselines, we observe that using a general-domain embedding model that has undergone supervised training (nomic-embed-text-v1super) as the starting checkpoint for domain adaptation, and further training it on biomedical data (nomic-embed-biosuper), results in lower performance than adapting a model pretrained only in an unsupervised manner on general-domain data (nomic-embed-text-v1unsup).
We further find that the adapter-based approach (E5-base-peft-biomed) degrades the performance of E5base, placing it behind other supervised baselines across all evaluated tasks. Additionally, tuning a biomedical language model (PubMedBERT) in a contrastive regime on biomedical data (PubMedBERT-embed) does not yield superior performance on any benchmark. This suggests that transferring contrastive knowledge from a strong general-domain embedding model is more effective than training a domain-specific LLM from scratch with a contrastive objective on in-domain data.
We also observe that competing baselines exhibit task-specific strengths: for example, all-mpnet-base-v2 performs particularly well on clustering tasks, while GTEbase and nomic-embed-text-v1 show strong performance in retrieval. In contrast, MOSAIC-Biosuper demonstrates consistently strong performance across task types, including semantic textual similarity (BIOSSES), clustering (MedrxivClusteringP2P), and retrieval (MedicalQARetrieval), achieving the highest average score overall.
However, as shown in Table 1, all of our models lag behind on the TRECCOVID dataset (Voorhees et al., 2021). Inspection of the TRECCOVID queries reveals that, alongside core biomedical and clinical questions, a significant fraction focuses on social or policy aspects of the pandemic (approximately 20%). Such queries, addressing societal impacts or interventions like school reopening, may fall outside the primary scope of the biomedical corpora used for model adaptation. This mismatch in domain coverage could partly explain the observed underperformance. Moreover, recent large-scale analyses of PubMed using embedding-based atlases have shown that COVID-19 literature forms a uniquely isolated cluster in embedding space, with strong internal topical fragmentation, further challenging biomedical models (Kobak et al., 2024). The qualitative analysis on BIOSSES is provided in the Appendix E.